OpenStudy (anonymous):

How many kilometers are there in 290,111 millimeters?

OpenStudy (ipwnbunnies):

There are 1000 millimeters in a meter. There are 1000 meters in a kilometer. \[290,111 mm * \frac{1 m}{1000 mm} * \frac{1 km}{1000 m} = ?\]

OpenStudy (somy):

Km*1000= m m*1000= mm so mm/ 1000000= Km

OpenStudy (ipwnbunnies):

Dividing by 1000 is the same as moving the decimal to the left 3 times, to save you some time. :3

OpenStudy (ipwnbunnies):

SOMY ♥

OpenStudy (anonymous):

Thanks again somy

OpenStudy (somy):

Bunnies ❤️

OpenStudy (somy):

you are welcome :)))

OpenStudy (anonymous):

;) & thanks Bunnies